How can I covert any quadratic equations like 3(x^2-2x)=2x^2-8 or x^2-3x-40=48 into standard/general form (ax^2+bx+c=0)?I know there is four ways to
covert, the two ways I need help with are converting by subrtracting a number from both sides and other expand quadratic equation and collect like terms.
Can you give me two examples for each solution?

Answer by rfer(16322)

3(x^2-2x)=2x^2-8
3x^2-6x-2x^2+8=0
x^2-6x+8=0
x^2-3x-40=48
x^2-3x-40-48=0
x^2-3x-88=0
Nothing different than manipulating any other equation.